The​ intensity, I, of light received at a source varies inversely as the square of the​ distance, d, from the source. If the light intensity is 63 ​foot-candles at 16 ​feet, find the light intensity at 8 feet.
Answer by jorel1380(3719)

Since the intensity varies inversely as the square of the distance, then:
63=n/(16)²
n=63x256=16128
16128/(8²)=16128/64=252 ft.candles as the intensity of light at 8 feet. ☺☺☺☺